OpenStudy (idealist10):

Evaluate the double integral 4x^3 dA, where R is the region bounded by the graphs of y=(x-1)^2 and y=-x+3.

OpenStudy (idealist10):

@ganeshie8 @phi @wio @radar

OpenStudy (idealist10):

Just tell me how to find the limits of integration for y. I can do rest of the problem. Because (x-1)^2=-x+3 so x=2, -1 for dx. But what about y?

OpenStudy (phi):

you could graph the two curves: a parabola and a line, using for example Geogebra

OpenStudy (idealist10):

But what are the limits of integration for y?

OpenStudy (phi):

I would make the inner integral over y the upper limit would be the blue line, and the lower limit the parabola these would be functions of x then the outer integral would be over x, with numeric limits

OpenStudy (idealist10):

So the limits of integration for y are 1 to 4? But the answer to this problem is 72/5.

OpenStudy (phi):

no, the limits for y are functions of x . in other words, depending on what x we are at, the y limits change (which we can see by looking at the graph) \[ \int_{-1}^2 \int_{(x-1)^2}^{3-x} 4 x^3 dy\ dx \]

OpenStudy (idealist10):

Thank you! Now I understand.